OpenStudy (anonymous):

State whether the relationship between the variables in the table is a direct variation, an inverse variation, or neither. If it is a direct or inverse variation, write a function to model it. x 3 9 13 20 y 9 27 39 60

OpenStudy (qwertty123):

check : if the product xy is constant for all pairs of x,y then the variation is inverse, but if the ratio y/x is constant for all pairs of x,y then the variation is direct.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

its direct

OpenStudy (qwertty123):

This is a direct variation, because as x increases, so does y. In other words both increase in the same direction, namely up. y = 3x

OpenStudy (anonymous):

y/x = 9/3 = 3 y/x = 27/9 =3 y/x = 39/13 =3 ?

OpenStudy (qwertty123):

YEAH You forgot the variable though!

OpenStudy (anonymous):

3x?

OpenStudy (anonymous):

9=k *3 ---> k=3 so, y=3x :)?

OpenStudy (qwertty123):

YEAH! :D

OpenStudy (anonymous):

woohoo!

OpenStudy (qwertty123):

YEAH! WAY TO GO! :D

OpenStudy (anonymous):

Thanks dude! :P

OpenStudy (qwertty123):

You welcome!